# Signature matrices of membranes

### v1

@@ Abstract
We prove that, unlike in the case of paths, the signature matrix of a membrane does not satisfy any algebraic relations. We derive novel closed-form expressions for the signatures of polynomial membranes and piecewise bilinear interpolations for arbitrary $2$-parameter data in $d$-dimensional space. We show that these two families of membranes admit the same set of signature matrices and scrutinize the corresponding affine variety.
@@ 1 Introduction
The signature of a path.
A path is a continuous map $X:[0,1]\to\mathbb{R}^{d}$ with piecewise continuously differentiable coordinate functions. Its $k$-th level signature $\sigma^{(k)}(X)$, as introduced by Chen in the seminal work [8], is a $k$-tensor of format $d\times\dots\times d$, defined by forming certain iterated integrals. In particular, its second level signature can be viewed as a signature matrix. It is defined as follows:
$\sigma^{(2)}(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{t_{2}}\partial X_{i}(t_{1})\,\partial X_{j}(t_{2})\;\mathrm{d}t_{1}\mathrm{d}t_{2}$ (1)
These matrices were studied in [2]. They satisfy some interesting properties: for example, their symmetric part has always rank $1$, due to the shuffle relations satisfied by path signatures.
Moreover, the set of signature matrices of polynomial paths with degree $\leq m$ and the set of signature matrices of piecewise linear paths with at most $m$ segments agree. The equations that vanish on these sets, that is, the varieties of signature matrices, are well-understood: they can be described as certain determinantal varieties [2, Theorem 3.4].
The signature of a membrane.
A membrane is a continuous map $X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}$ where the coordinate functions $X_{1},X_{2},\dots,X_{d}$ are (piecewise) continuously differentiable functions in the two arguments. In particular the differentials $\partial_{12}X_{j}(s,t):=\frac{\partial^{2}}{\partial s\partial t}X_{j}(s,t)$ obey the usual rules of calculus. Generalizing the path setting, one can associate a collection of tensors to $X$ via iterated integrals (see Definition 2.2). In particular, the signature matrix of a membrane $X$ is defined as an $\mathbb{R}$-valued $d\times d$-matrix with entries
$\sigma^{(2)}(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{1}\int_{0}^{t_{2}}\int_{0}^{s_{2}}\partial_{12}X_{i}(s_{1},t_{1})\,\partial_{12}X_{j}(s_{2},t_{2})\;\mathrm{d}s_{1}\mathrm{d}t_{1}\mathrm{d}s_{2}\mathrm{d}t_{2}$ (2)
for every $1\leq i,j\leq d$. This matrix has recently been introduced as the second level of the two-parameter $\mathsf{id}$-signature [26, 48, 14].
While it is easy to see that, contrary to path signatures, this signature does not satisfy shuffle relations, little is known about the algebraic relations satisfied by signature tensors in general. In this paper, inspired by [2], we address this question for the case of matrices. We write $\sigma(X):=\sigma^{(2)}(X)$ in the following. We focus on polynomial membranes of order $(m,n)$, that is, membranes $X$ whose component functions are polynomials of multidegree $\leq(m,n)$, and piecewise bilinear membranes of order $(m,n)$, see Definition 3.14. Their signature matrices (2) form semialgebraic subsets of $\mathbb{R}^{d^{2}}$. To study the polynomials that vanish on these sets, we move to the algebraically closed field $\mathbb{C}$ and consider the Zariski closures
$\mathcal{P}_{d,m,n}^{\mathsf{pol}}\quad\text{ and }\quad\mathcal{P}_{d,m,n}^{\mathsf{bil}}.$ (3)
Main question.
What can be said about the varieties of signature matrices of polynomial membranes and piecewise bilinear membranes?
@@
An implicitization problem.
Our strategy is to view the varieties (3) as solutions to an implicitization problem. To this end, we show that both classes of membranes admit a dictionary membrane (see [41]), i.e. a membrane which yields all other membranes in the class via linear transforms, up to certain terms that vanish under the signature map.
Our first candidate of a dictionary is the $mn$-dimensional moment membrane,
@@
where $\otimes$ is the Kronecker product. If $m=n=2$, then $\mathsf{Mom}^{2,2}(s,t)=(st,st^{2},s^{2}t,s^{2}t^{2})$ has the signature matrix
$\sigma(\mathsf{Mom}^{2,2})=\begin{pmatrix}\frac{1}{4}&\frac{1}{3}&\frac{1}{3}&\frac{4}{9}\\[4.0pt] \frac{1}{6}&\frac{1}{4}&\frac{2}{9}&\frac{1}{3}\\[4.0pt] \frac{1}{6}&\frac{2}{9}&\frac{1}{4}&\frac{1}{3}\\[4.0pt] \frac{1}{9}&\frac{1}{6}&\frac{1}{6}&\frac{1}{4}\end{pmatrix}=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}$ (5)
with a Kronecker factorization into path signatures (Corollary 3.6). Note that indeed every $d$-dimensional polynomial membrane $X$ with terms of multidegree $\geq(1,1)$ is of the form $X=A\,\mathsf{Mom}^{m,n}$. With equivariance of the signature (Lemma 2.6) we obtain the closed formula,
$\sigma(X)=A\sigma(\mathsf{Mom}^{m,n})A^{\top},$ (6)
as a matrix congruence. Hence, continuing with our example (5) and $d=2$, we have
@@
Thus, we obtain the first entry of our signature matrix from (5) and (6): It is the homogeneous polynomial
$\sigma(X)_{1,1}=\frac{1}{4}a_{1,1}^{2}$ $+\frac{1}{2}a_{1,1}a_{1,2}+\frac{1}{2}a_{1,1}a_{1,3}+\frac{5}{9}a_{1,1}a_{1,4}+\frac{1}{4}a_{1,2}^{2}$
@@
of degree two. The remaining three polynomial entries of $\sigma(X)$ are provided in Example 3.7. We illustrate the polynomial membrane (7) for explicit $A$ in Figure 3.
The axis membrane.
Interpolation is a necessary step to fit discrete data into the framework of iterated integrals. Our second important dictionary, the axis membrane $\mathsf{Axis}^{m,n}$, is inspired by the dictionary for piecewise linear paths [2, 41] (see Definition 3.15). This $mn$-dimensional membrane allows us to interpolate arbitrary $d$-dimensional discrete data on an $m\times n$ grid via a piecewise bilinear membrane. For an illustration of such an interpolation see $Y$ from Figure 3. We provide novel closed-form expressions and computational complexity statements for its signatures (Theorem 3.19).
Main results.
Using our description via dictionaries, we obtain an analogue of [2, Theorem 3.3] for membranes:
Theorem (4.1).
For all $d,m$ and $n$,
$\mathcal{P}^{\mathsf{pol}}_{d,m,n}=\mathcal{P}^{\mathsf{bil}}_{d,m,n}=:\mathcal{P}_{d,m,n}.$
In fact, already the defining semi-algebraic sets agree.
These varieties are irreducible (Proposition 4.3) and we obtain a grid
@@
$\mathcal{P}_{d,1,n}\subseteq\dots\subseteq\mathcal{P}_{d,M-1,n}\subseteq\mathcal{P}_{d,M,n}=\mathcal{P}_{d,M+1,n}$ =$\dots$= …
for all $m,n\in\mathbb{N}$. As $\mathcal{P}_{d,1,m}=\mathcal{P}_{d,m,1}$ agrees with the path varieties studied in [2] (Lemma 4.5), it follows that $\mathcal{P}_{d,1,N}=\mathcal{P}_{d,M,1}$ is the universal path variety. However, in strong contrast to the path case, in general there are no algebraic relations in our two-parameter signature matrices (2) at all:
@@
The limit variety $\mathcal{P}_{d,M,N}$ in (8) attains the full ambient dimension, i.e. $\mathcal{P}_{d,M,N}=\mathbb{C}^{d^{2}}$.
The varieties become more interesting if $m$ and $n$ are small compared to $d$. Employing the theory of orbits under congruence [45, 30] we were able to find explicit formulas for the dimensions of $\mathcal{P}_{d,m,n}$ whenever $mn\leq d$, generalizing [2, Theorem 3.4]. Note that there are $dmn\leq d^{2}$ parameters involved, so the dimension can be at most $dmn$.
@@
For $mn\leq d$ the dimension of $\mathcal{P}_{d,m,n}$ is
These results are complemented by some explicit computations using the open-source computer algebra systems Macaulay 2 [28, 19], and OSCAR [40, 13], see e.g. Examples 4.12 and 4.2, Remark 4.13 and Example 4.14.
@@
The collection of iterated integrals of a path has been used as a tool for classification in various settings since Chen’s seminal 1954 work [8]. After pioneering work by Lyons and collaborators on the use of signatures as features for data streams in the early 2010s [29, 32], there has been a surge in the number of applications, which range from motion recognition [47], finance [4], deep learning [33, 34, 46, 35], oceanography [44], and lately data science for images [48, 31]. We refer the interested reader to the surveys [10, 38].
Since [2, 41, 24] established the initial connection between stochastic analysis and algebraic geometry, the signature has also been applied in various related fields, including the theory of tensor decompositions [3, 22, 25], clustering tasks [11, 39], tropical time series analysis [15, 16] and the geometry of paths [17, 1, 42].
Recently, the signature has been extended to two parameters, generalizing from paths to membranes: the mapping space signature [26] provides a principled two-parameter extension of Chen’s iterated integrals, based on its original use in topology, whereas the two-parameter sums signature [18] characterizes equivalence up to warping in two directions. These two approaches have recently been combined in [14] using mixed partial derivatives. Due to the lack of a two-parameter version of Chen’s identity in all the above references, recent works [9, 37, 36] use higher category theory to find an appropriate generalization of a group that is rich enough to represent the horizontal and vertical concatenation of membranes.
@@ 2 Iterated integrals
The iterated-integrals signature $\sigma=\sigma(X)$ of a $d$-dimensional path $X:[0,1]\to\mathbb{R}^{d}$ is an element of the tensor series algebra $T((\mathbb{R}^{d}))\cong\operatorname{Hom}(T(\mathbb{R}^{d}),\mathbb{R})$, defined by
$\sigma:T(\mathbb{R}^{d})$ $\to\mathbb{R},$
$e_{i_{1}}\otimes\dots\otimes e_{i_{k}}$ $\mapsto\int_{\Delta}\mathrm{d}X_{i_{1}}(t_{1})\dots\mathrm{d}X_{i_{k}}(t_{k}),$ (9)
where $\Delta$ is the simplex $10<t_{1}<\dots<t_{k}<$ 1. It is common to write the tensor $e_{i_{1}}\otimes\dots\otimes e_{i_{k}}$ as a word $i_{1}\dots i_{k}$, identifying the tensor algebra with the free associative algebra over the letters 1 to d. However, the tensor algebra viewpoint allows us to consider the signature for paths (and later membranes) in any $\mathbb{R}$-vector space $V$, without choosing a basis, which will be useful in the following. Let us first recall some examples in the (classical, one-parameter) path setting.
@@
    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the linear path $[0,1]\rightarrow\mathbb{R}^{d},t\mapsto ut$ is given by
    $\sigma^{(k)}_{i_{1},\dots,i_{k}}=\frac{u_{i_{1}}\dots u_{i_{k}}}{k!}$ (10)
@@
    Let $\mathsf{Axis}^{d}:[0,1]\to\mathbb{R}^{d}$ be the canonical axis path from [2, Example 2.1] or [41, eq. (8)] of order $d$, i.e., the unique piecewise linear path with $d$ segments whose graph interpolates the $d+1$ equidistant support points $(\frac{i}{d},\sum_{1\leq j\leq i}e_{j})$ with $0\leq i\leq d$. This path is illustrated in Figure 2 for $d=3$. For all $d$, its signature matrix is the upper triangular matrix
    $\sigma^{(2)}_{i_{1},i_{2}}=\begin{cases}1&\text{if }i_{1}<i_{2}\\ \frac{1}{2}&\text{if }i_{1}=i_{2}\\ 0&\text{if }i_{1}>i_{2}.\end{cases}$ (11)
@@
    $\sigma^{(k)}_{i_{1},i_{2},\dots,i_{k}}=\frac{i_{2}i_{3}\dots i_{k}}{(i_{1}+i_{2})(i_{1}+i_{2}+i_{3})\dots(i_{1}+\dots+i_{k})}$ (12)
    for all $i_{1},\dots,i_{k}$ due to [2, Example 2.2].
@@
In the following we recall the ‘nearest’ extension of the path signature (2) for multiple parameters and comment on its various modifications in Remark 2.3.
Definition 2.2.
Let $V$ be a finite dimensional vector space and $X:[0,1]^{p}\to V$ a Lipschitz continuous map. Let $\partial_{1\dots p}:=\frac{\partial^{p}}{\partial x_{1}\dots\partial x_{p}}$ be the linear partial differential operator. Then, since both the integral and $\partial_{1\dots p}$ are linear operators, the following linear form $\sigma=\sigma(X)\in T(V^{*})^{*}$ is well-defined:
$\sigma:T(V^{*})$ $\to\mathbb{R},$
$\alpha_{1}\otimes\dots\otimes\alpha_{k}$ $\mapsto\int_{\Delta^{p}}\partial_{1\dots p}(\alpha_{1}\circ X)(t_{11},\dots,t_{1p})\dots\partial_{1\dots p}(\alpha_{k}\circ X)(t_{k1},\dots,t_{kp})\,\mathrm{d}\mathbf{t}$
where $\mathrm{d}\mathbf{t}=\mathrm{d}t_{11}\dots\mathrm{d}t_{1p}\dots\mathrm{d}t_{k1}\dots\mathrm{d}t_{kp}$ and $\Delta^{p}=\Delta\times\dots\times\Delta$ is the Cartesian product of $p$ simplices. We call $\sigma$ the iterated-integrals signature of $X$. The $k$-th level signature $\sigma^{(k)}(X)$ is the restriction of $\sigma(X)$ along $(V^{*})^{\otimes k}\to T(V^{*})$, in particular it is naturally an element of $V^{\otimes k}$.
@@
    For the one-parameter framework $p=1$ in $d$-dimensional space $V=\mathbb{R}^{d}$ equipped with standard basis $e_{1},\dots,e_{d}$ we obtain the classical formulation (2) for paths. We refer the interested reader to the text books [23, 21].
@@
    When ‘replacing’ our square increments $\partial_{1\dots p}$ with Jacobian minors we obtain the $\mathsf{id}$-signature introduced in [26].

### v2

@@ Abstract
The signature of a membrane is a sequence of tensors whose entries are iterated integrals. We study algebraic properties of membrane signatures, with a focus on signature matrices of polynomial and piecewise bilinear membranes. Generalizing known results for path signatures, we show that the two families of membranes admit the same set of signature matrices and we examine the corresponding affine varieties. In particular, we prove that there are no algebraic relations on signature matrices of membranes, in contrast to the path case. We complement our results by a linear time algorithm for the computation of signature tensors for piecewise bilinear membranes.
@@ 1 Introduction
A path is a continuous map $X:[0,1]\to\mathbb{R}^{d}$ with piecewise continuously differentiable coordinate functions. The signature of a path was introduced by Chen in his seminal work [Chen1954IteratedIA] and has since become a central construction in the theory of rough paths. The second level signature $S(X)\in\mathbb{R}^{d\times d}$ of a path $X$ is a matrix whose entries are iterated integrals
$S(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{t_{2}}\dot{X}_{i}(t_{1})\,\dot{X}_{j}(t_{2})\;\mathrm{d}t_{1}\mathrm{d}t_{2}$
corresponding to $1\leq i,j\leq d$, and where $\dot{X}$ denotes the derivative of $X$. These signature matrices were studied in [bib:AFS2019, Section 3] and satisfy several interesting properties: their symmetric parts always have rank $1$, due to the shuffle relations satisfied by path signatures. The sets of signature matrices of polynomial and piecewise linear paths coincide. This set is semi-algebraic and its Zariski closure can be described as a certain determinantal variety.
The signature of a membrane
Recently, the signature has been extended to membranes. A $d$-dimensional membrane is a continuous map $X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}$ where the coordinate functions $X_{1},X_{2},\dots,X_{d}$ are (piecewise) continuously differentiable functions in the two arguments. In particular the differentials $\partial_{12}X_{j}(s,t):=\frac{\partial^{2}}{\partial s\partial t}X_{j}(s,t)$ obey the usual rules of calculus. Generalizing the path setting, one can associate a collection of tensors to $X$ via iterated integrals (see Definition 2.2). The associated object was termed $\mathsf{id}$-signature in [diehl2024signature, Definition 4.1], [GLNO2022, Section 4.7]. We call the second level of the $\mathsf{id}$-signature the signature matrix of a membrane. It is defined as the $\mathbb{R}$-valued $d\times d$ matrix $S(X)$ with entries
$S(X)_{i,j}$ $:=\int_{0}^{1}\int_{0}^{1}\int_{0}^{t_{2}}\int_{0}^{s_{2}}\partial_{12}X_{i}(s_{1},t_{1})\,\partial_{12}X_{j}(s_{2},t_{2})\;\mathrm{d}s_{1}\mathrm{d}t_{1}\mathrm{d}s_{2}\mathrm{d}t_{2}$ (1)
for every $1\leq i,j\leq d$.
Although it is easy to see that the $\mathsf{id}$-signature does not satisfy the same algebraic relations as the path signature [diehl2024signature, Section 4.3], little is known about its algebraic relations in general. In this paper, inspired by [bib:AFS2019], we address this question for the case of signature matrices. For this we consider the family of membranes whose component functions are polynomials of order $(m,n)$. Their signature matrices (1) form a semialgebraic subset of $\mathbb{R}^{d\times d}$. To study the polynomials that vanish on these sets, we move to the algebraically closed field $\mathbb{C}$ and consider the associated Zariski closure, in the following denoted by $\mathcal{M}_{d,m,n}$.
@@
An implicitization problem
Our strategy is to view the variety $\mathcal{M}_{d,m,n}$ as the Zariski closure of a certain matrix congruence orbit. For this we use a dictionary membrane inspired by [bib:PSS2019], i.e., a special membrane that yields all polynomial membranes via a suitable linear transform, up to terms that vanish under the signature map. The moment membrane,
@@
is a dictionary for polynomial membranes of order $(m,n)$, where $\otimes$ denotes the Kronecker product. We call this membrane a product membrane since it factorizes into paths. We will see in Section 3 that the signature matrix of a product membrane is a Kronecker product of the signature matrices of its path factors. If for example $m=n=2$, then $\mathsf{Mom}^{2,2}(s,t)=(st,st^{2},s^{2}t,s^{2}t^{2})$ has the signature matrix
$S(\mathsf{Mom}^{2,2})=\begin{pmatrix}\frac{1}{4}&\frac{1}{3}&\frac{1}{3}&\frac{4}{9}\\[4.0pt] \frac{1}{6}&\frac{1}{4}&\frac{2}{9}&\frac{1}{3}\\[4.0pt] \frac{1}{6}&\frac{2}{9}&\frac{1}{4}&\frac{1}{3}\\[4.0pt] \frac{1}{9}&\frac{1}{6}&\frac{1}{6}&\frac{1}{4}\end{pmatrix}=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}.$ (3)
Since the moment membrane is a dictionary, we can express every polynomial membrane $X$ without linear terms of order $(m,n)$ as a linear transform $X=A\,\mathsf{Mom}^{m,n}$, where the matrix $A$ is chosen accordingly. Note that the membrane signature is invariant under addition of linear terms. With equivariance of the signature (Lemma 2.6) we obtain the closed formula,
$S(X)=AS(\mathsf{Mom}^{m,n})A^{\top},$ (4)
through a matrix congruence transform. Continuing with our example (3) and $d=2$, we have
@@
and thus with (3) and (4) the homogeneous polynomial
$S(X)_{1,1}=\frac{1}{4}a_{1,1}^{2}$ $+\frac{1}{2}a_{1,1}a_{1,2}+\frac{1}{2}a_{1,1}a_{1,3}+\frac{5}{9}a_{1,1}a_{1,4}+\frac{1}{4}a_{1,2}^{2}$
@@
as the first entry of our signature matrix. The remaining $3$ polynomial entries of $S(X)$ are provided in Example 3.7. We illustrate the polynomial membrane (5) for an explicit matrix $A$ in Figure 3.
Main results
The varieties of polynomial signature matrices $\mathcal{M}_{d,m,n}$ can be arranged in the following grid of inclusions,
@@
$\mathcal{M}_{d,1,n}\subseteq\dots\subseteq\mathcal{M}_{d,M-1,n}\subseteq\mathcal{M}_{d,M,n}=\mathcal{M}_{d,M+1,n}$ =$\dots$= …
for all $m,n\in\mathbb{N}$. We show in Lemma 4.5 that $\mathcal{M}_{d,1,m}=\mathcal{M}_{d,m,1}$ agrees with the path varieties studied in [bib:AFS2019]. In particular, it follows that $\mathcal{M}_{d,1,N}=\mathcal{M}_{d,M,1}$ is the variety parametrizing the signature matrices of all smooth paths, called universal path variety in [bib:AFS2019, Section 4.3]. However, in sharp contrast to the path case, there are generally no algebraic relations in our two-parameter signature matrices (1) at all.
@@
The variety $\mathcal{M}_{d,M,N}$ in (6) attains the full ambient dimension, i.e. $\mathcal{M}_{d,M,N}=\mathbb{C}^{d\times d}$. More precisely, $\mathcal{M}_{d,m,n}=\mathbb{C}^{d\times d}$ whenever $m+n>d$ with $m,n\not=1$.
The proof relies on the theory of matrix congruence orbits [TERAN201144, HORN20061010]. Using this machinery, we can find lower and upper bounds for our varieties $\mathcal{M}_{d,m,n}$:
@@
Let $\mathcal{S}_{a,b}$ be the variety of $d\times d$ matrices whose symmetric part has rank $\leq a$ and whose skew-symmetric part has rank $\leq b$. Then
  • $\mathcal{S}_{(m-2)(n-2),m+n-2}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{mn,m+n-2}$ if $m,n$ are even,
  • $\mathcal{S}_{2,n-1}\subseteq\mathcal{M}_{d,2,n}\subseteq\mathcal{S}_{2(n-1)+1,n+1}$ if $n$ is odd,
  • $\mathcal{S}_{(m-2)(n-1)-1,m+n-1}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{m(n-1)+1,m+n-1}$ if $m\geq 4$ is even, $n$ is odd,
  • $\mathcal{S}_{(m-1)(n-1)+1,m+n-2}=\mathcal{M}_{d,m,n}$ if $m,n$ are odd.
In particular, it follows that our varieties are only interesting if $m$ and $n$ are small compared to $d$. From [TERAN201144] we obtain explicit formulas for the dimensions of $\mathcal{M}_{d,m,n}$ whenever $mn\leq d$, generalising [bib:AFS2019, Theorem 3.4]. Note that there are $dmn\leq d^{2}$ parameters involved, so the dimension can be at most $dmn$.
Theorem (4.7).
For $mn\leq d$, the dimension of $\mathcal{M}_{d,m,n}$ is
  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+4(m+n)-4$, if $m$ and $n$ even,
  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{3}{2}mn+m+n$, if $m$ even and $n$ odd,
  • $dmn-\frac{1}{2}m^{2}n^{2}+m^{2}(n-1)+(m-1)n^{2}-\frac{7}{2}mn+3(m+n)-2$, if $m$ and $n$ odd.
The results are complemented by some explicit computations using the open-source computer algebra system Macaulay 2 [GraysonStillmanMacaulay2, eisenbud2001computations], see e.g. Examples 4.12 and 4.2, Remark 4.14 and Example 4.15.
Piecewise bilinear membrances
Bilinear interpolation is an easy way to fit discrete data into the framework of iterated integrals. It allows us to interpolate arbitrary $d$-dimensional discrete data on an $m\times n$ grid via a piecewise bilinear membrane of order $(m,n)$; see Definition 3.13. For an illustration of such an interpolation see $Y$ from Figure 3. In Definition 3.14 we define our second important dictionary, the axis membrane $\mathsf{Axis}^{m,n}$. It is inspired by the dictionary for piecewise linear paths from [bib:PSS2019], and it allows an analogue of [bib:AFS2019, Theorem 3.3] for membranes.
Theorem (4.1).
The sets of signature matrices of polynomial and piecewise bilinear membranes with order $(m,n)$ coincide. In particular, both sets define the same algebraic variety $\mathcal{M}_{d,m,n}$.
A naive computation of signature matrices via congruence operation results only in a quadratic time algorithm with respect to the number of interpolation points; see Corollary 3.11. In particular, there is no analogue of Chen’s identity for the id-signature [GLNO2022, diehl2024signature]. However, using a cumulated approach inspired by [DEFT22, DS22] we can drop this cost to linear complexity.
Corollary (3.19).
The signature matrix of a piecewise bilinear membrane in $\mathbb{R}^{d}$ of order $(m,n)$ is computable in $\mathcal{O}(d^{2}mn)$.
We can generalize this idea for the $k$-th level of the $\mathsf{id}$-signature tensor; see Theorem 3.18. This yields a linear two-parameter integration algorithm for the signatures of images.
The code accompanying this paper is available at
https://github.com/felixlotter/signature-matrices-of-membranes.
@@
The collection of iterated integrals of a path has been used as a tool for classification in various settings since Chen’s seminal 1954 work [Chen1954IteratedIA]. After pioneering work by Lyons and collaborators on the use of signatures as features for data streams in the early 2010s [Gyu13, handw13], there has been a surge in the number of applications.
Since [bib:AFS2019, bib:PSS2019, GALUPPI2019282] established the initial connection between stochastic analysis and algebraic geometry, the signature has also been applied in various related fields, including the theory of tensor decompositions [AGOSS2025, FLS24, galuppi2024ranksymmetriessignaturetensors], clustering tasks [clausel2024barycenterfreenilpotentlie, mignot:hal-04392568], tropical time series analysis [DEFT22, diehl2023fruitsfeatureextractionusing] and the geometry of paths [Diehl2019, 10.1007/978-3-031-38271-0_45, preiß2024algebraicgeometrypathsiteratedintegrals].
Recently, the signature has been extended to two parameters, generalizing from paths to membranes: the mapping space signature [GLNO2022] provides a principled two-parameter extension of Chen’s iterated integrals, based on its original use in topology, whereas the two-parameter sums signature [DS22] characterizes equivalence up to warping in two directions. These two approaches have recently been combined in [diehl2024signature] using mixed partial derivatives. Due to the lack of a two-parameter version of Chen’s identity in all the above references, recent works [CDEFT2024, LO2023, lee2024] use higher category theory to find an appropriate generalization of a group that is rich enough to represent the horizontal and vertical concatenation of membranes.
Outline.
In section Section 2 we recall the notion of iterated-integrals signatures for paths and membranes. Section 3 introduces the notion of a product membrane. For arbitrary $k$, the $k$-th level signature of these membranes can be computed from path signatures, which will allow us to use theory from the path setting. We define the moment membrane and the canonical axis membrane and include algorithmic considerations of bilinear interpolations and linearly computable signature entries in Section 3.2. Finally, we return to the matrix case $k=2$ in Section 4 where we prove our main results on signature matrix varieties. For open problems and future work see Section 5.
@@ 2 Iterated integrals
Let $V:=\mathbb{R}^{d}$ and let $T(V^{*})$ denote the tensor algebra $\bigoplus_{k\in\mathbb{N}_{0}}(V^{*})^{\otimes k}$. The iterated-integrals signature $\sigma(X)$ of a $d$-dimensional path $X:[0,1]\to V$ is an element of the tensor series algebra $T((V^{*}))\cong\operatorname{Hom}(T(V),\mathbb{R})$, defined by
$\sigma(X):T(V^{*})$ $\to\mathbb{R},$
$e_{i_{1}}^{*}\otimes\dots\otimes e_{i_{k}}^{*}$ $\mapsto\int_{\Delta_{k}}\mathrm{d}X_{i_{1}}(t_{1})\dots\mathrm{d}X_{i_{k}}(t_{k}),$ (7)
where $\Delta_{k}$ is the simplex $0\leq t_{1}\leq\dots\leq t_{k}\leq 1$. The signature matrix of $X$ is $S_{ij}(X):=\sigma(X)(e_{i}^{*}\otimes e_{j}^{*})$. It is common to write the tensor $e^{*}_{i_{1}}\otimes\dots\otimes e^{*}_{i_{k}}$ as a word $\mathtt{i}_{1}\dots\mathtt{i}_{k}$, identifying the tensor algebra $T(V^{*})$ with the free associative algebra $\mathbb{R}\langle\mathtt{1},\ldots,\mathtt{d}\rangle$ over the letters 1 to d. However, the tensor algebra viewpoint allows us to consider the signature for paths (and later membranes) in any $\mathbb{R}$-vector space $V$, without choosing a basis, which will be useful in the following. Let us first recall some examples in the (classical, one-parameter) path setting.
@@
    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the linear path $X:[0,1]\rightarrow\mathbb{R}^{d},t\mapsto ut$ is given by
    $\sigma(X)(e_{i_{1}}^{*}\otimes\dots\otimes e_{i_{k}}^{*})=\frac{u_{i_{1}}\dots u_{i_{k}}}{k!}$ (8)
@@
    Let $\mathsf{Axis}^{d}:[0,1]\to\mathbb{R}^{d}$ be the canonical axis path from [bib:AFS2019, Example 2.1] or [bib:PSS2019, eq. (8)] of order $d$, i.e., the unique piecewise linear path with $d$ segments whose graph interpolates the $d+1$ equidistant support points $(\frac{i}{d},\sum_{1\leq j\leq i}e_{j})$ with $0\leq i\leq d$. This path is illustrated in Figure 2 for $d=3$. For all $d$, its signature matrix is the upper triangular matrix
    $S_{i_{1},i_{2}}(\mathsf{Axis}^{d})=\begin{cases}1&\text{if }i_{1}<i_{2}\\ \frac{1}{2}&\text{if }i_{1}=i_{2}\\ 0&\text{if }i_{1}>i_{2}.\end{cases}$ (9)
@@
    $\sigma(\mathsf{Mom}^{d})(e_{i_{1}}^{*}\otimes\dots\otimes e_{i_{k}}^{*})=\frac{i_{2}i_{3}\dots i_{k}}{(i_{1}+i_{2})(i_{1}+i_{2}+i_{3})\dots(i_{1}+\dots+i_{k})}$ (10)
    for all $i_{1},\dots,i_{k}$ due to [bib:AFS2019, Example 2.2].
@@
The path signature (2) can be extended to two parameters in multiple ways. We will study the $\mathsf{id}$-signature as defined in [diehl2024signature, Definition 4.1]. We comment on its various modifications in Remark 2.3.
Let $V$ be a finite dimensional $\mathbb{R}$-vector space. A membrane in $V$ is a Lipschitz continuous map $X:[0,1]^{2}\to V$.
Definition 2.2 ($\mathsf{id}$-signature).
Let $V$ be a finite dimensional vector space and $X$ a membrane in $V$. Let $\partial_{12}:=\frac{\partial^{2}}{\partial x_{1}\partial x_{2}}$ be the linear partial differential operator. Then, since both the integral and $\partial_{12}$ are linear operators, the following linear form $\sigma(X)\in T(V^{*})^{*}$ is well-defined:
$\sigma(X):T(V^{*})$ $\to\mathbb{R},$
$\alpha_{1}\otimes\dots\otimes\alpha_{k}$ $\mapsto\int_{(\mathbf{s},\mathbf{t})\in\Delta^{2}_{k}}\partial_{12}(\alpha_{1}\circ X)(s_{1},t_{1})\dots\partial_{12}(\alpha_{k}\circ X)(s_{k},t_{k})\,\mathrm{d}\mathbf{s}\mathrm{d}\mathbf{t}$
where $\mathrm{d}\mathbf{s}=\mathrm{d}s_{1}\dots\mathrm{d}s_{k}$, $\mathrm{d}\mathbf{t}=\mathrm{d}t_{1}\dots\mathrm{d}t_{k}$ and $\Delta_{k}^{2}=\Delta_{k}\times\Delta_{k}$. We call $\sigma(X)$ the $\mathsf{id}$-signature of $X$. The $k$-th level signature $\sigma^{(k)}(X)$ is the restriction of $\sigma(X)$ along $(V^{*})^{\otimes k}\to T(V^{*})$, in particular it is naturally an element of $V^{\otimes k}\cong((V^{*})^{\otimes k})^{*}$.
@@
    When ‘replacing’ our square increments $\partial_{12}$ with Jacobian minors we obtain the $\mathsf{id}$-signature introduced in [GLNO2022, Section 4.7].
@@
    For membranes and standard basis $e_{1}^{*},\dots,e_{d}^{*}$ of $V^{*}=(\mathbb{R}^{d})^{*}$, we obtain the $\mathsf{id}$-signature according to [diehl2024signature]. Its level $k$ component can be stored in the $d\times\dots\times d$ array,
    $\sigma^{(k)}_{i_{1},\dots,i_{k}}(X)$ $:=\sigma(e_{i_{1}}^{*}\otimes\dots\otimes e_{i_{k}}^{*})$
    $=\int\limits_{\begin{subarray}{c}0\leq s_{1}\leq\dots\leq s_{k}\leq 1\\ 0\leq t_{1}\leq\dots\leq t_{k}\leq 1\end{subarray}}\partial_{12}X_{i_{1}}(s_{1},t_{1})\dots\partial_{12}X_{i_{k}}(s_{k},t_{k})\,\mathrm{d}s_{1}\mathrm{d}t_{1}\dots\mathrm{d}s_{k}\mathrm{d}t_{k}.$
    In particular, for level $k=2$, we obtain our signature matrix $S(X)=\sigma^{(2)}(X)$ from the introduction, (1).